OpenStudy (anonymous):

Quadratic Formula x²=-2x+33 ???

OpenStudy (anonymous):

x^2 + 2x -33=0 a=1 b=2 c=-33 x=(-b+(sqrt(b^2-4ac)))/2a or x=(-b-(sqrt(b^2-4ac)))/2a now subsitute for a and b and c

OpenStudy (anonymous):

help :/

OpenStudy (anonymous):

I told you what to do well do it :)

OpenStudy (anonymous):

_________ -b + /b^2 - 4ac ---------------- : ) 2a im just re word ing

OpenStudy (anonymous):

but can someone tell me what to do next . ? please

OpenStudy (anonymous):

substitute the values a b c

OpenStudy (anonymous):

I understand that but hmmm..

OpenStudy (anonymous):

a = x^2 b = x c = the constant

OpenStudy (anonymous):

if a=10 and b=5 a+b=15 similarly x=(2+sqrt(4+4*33))/2=1+sqrt(34)

OpenStudy (anonymous):

-b + <----- that plus its a (plus or minus) just symbol on chart

OpenStudy (anonymous):

so u get 2 answers

OpenStudy (anonymous):

okay I understand so far and then what

OpenStudy (anonymous):

What do you want to after finding the answer??? :D lol